OpenStudy (anonymous):

someone help me please ... 9x^2+4x+5=0 solve

Parth (parthkohli):

Okay. What method are you comfortable with? Factoring? Quadratic formula?

OpenStudy (anonymous):

whatever ... just solve it ...

Parth (parthkohli):

Though, I'd suggest the formula here.\[x = {-b \pm \sqrt{b^2 - 4ac} \over 2a} \]

Parth (parthkohli):

\[\implies x = {-4 \pm \sqrt{4^2 - 4(9)(5)} \over 2(9)} \]Can you solve it?

OpenStudy (angela210793):

does lovebird know wht a,b,c are? O.o

Parth (parthkohli):

Okay.. wait.. \(9x^2 + 4x + 5 = 0\) a b c

Parth (parthkohli):

\(a\) is the coefficient of \(x^2\). \(b\) is the coefficient of \(x^1\) \(c\) is the coefficient of \(x^0\) or \(1\).

OpenStudy (anonymous):

oh then what can i doo @ParthKohli : i like you ... thnx :D

Parth (parthkohli):

Okay, so can you continue with that? Thanks!

OpenStudy (anonymous):

noo i cant .... help me

OpenStudy (anonymous):

i suck at maths

Parth (parthkohli):

Okay.\[x = {-4\pm \sqrt{4^2 - 4(9)(5)} \over 2(9)} \]\[x = {-4 \pm \sqrt{16 - 180} \over 18} \]\[x = {-4 \pm \sqrt{-164} \over 18} \]Before continuing, do you know what \(\pm\) is?

OpenStudy (anonymous):

yes i know that